value: 5.00 points Shannon Company seg ents its income statement in o is North and South Divisions. T e companys ove all sales con bu on margin at , and ne opera The North Divisions contribution margin and contribution margin ratio are S 165,000 and 75% respective The South Division s segment margin is 46 that cannot be traced to either division. g in ome are S5 The company has 00 56, 60%, an S5600 re e s of common fixed expenses Required: Prepare an income statement for Shannon Company that uses the contr bution format and is segmented by divisions. (Do not round your intermediate percentage answers and round your final percentage answers to 1 decimal place (i.e. .1234 should be entered as 12.3).) Divisions Total Company North South Amount Amount Amount
Answer #1

Income Statement

Divisions

Total Company

North

South

Amount

%

Amount

%

Amount

%

Sales

$    5,60,000.00

100.0%

$   2,20,000.00

100%

$         3,40,000.00

100.0%

Variable cost

$    2,24,000.00

40.0%

$       55,000.00

25%

$         1,69,000.00

49.7%

Contribution margin

$    3,36,000.00

60.0%

$   1,65,000.00

75%

$         1,71,000.00

50.3%

Division Fixed cost

$    2,24,000.00

40.0%

$       99,000.00

45%

$         1,25,000.00

36.8%

Segment Margin

$    1,12,000.00

20.0%

$       66,000.00

30%

$             46,000.00

13.5%

Fixed cost

$        56,000.00

10.0%

Net income

$        56,000.00

10.0%

Working

Income Statement

Divisions

Total Company

North

South

Amount

Amount

Amount

Sales

560000

=165000/75%

=560000-220000

Variable cost

=560000-336000

=220000-165000

=340000-171000

Contribution margin

=560000*60%

165000

=336000-165000

Division Fixed cost

=99000+125000

=165000-66000

=171000-46000

Segment Margin

=56000+56000

=112000-46000

46000

Fixed cost

56000

Net income

56000